# NIMO: a Nonlinear Interpretable MOdel

### v1

@@ 5 Conclusions
In this paper we propose a nonlinear interpretable model. Different from other hybrid approaches, our model retains the interpretability of linear coefficients – just as in standard linear models – while also allowing for nonlinear interaction terms. However, unlike linear models, we are not limited to global linear coefficients that operate at the population level. Our model can also learn local nonlinearities, enabling the generation of modified linear coefficients tailored to individual data points. This can be viewed as an individual correction on top of the linear prediction. In order to effectively train the proposed approach, we propose an optimization algorithm based on profile likelihood and adaptive ridge regression. Specifically, we are able to optimize the neural network parameters via gradient descent while analytically updating the linear coefficients. By means of sparsity in the neural networks we can also automatically select relevant nonlinear interactions. We showcase the proposed approach on synthetic dataset and show that we can learn the underlying coefficients and nonlinearities, outperforming nonlinear and hybrid approaches. We also showcase the model on real world datasets and confirmed the same findings. When the dataset does not contain nonlinear interaction terms we recover sparse coefficients and the network becomes less relevant. When the dataset contains nonlinearities we can learn them with the same accuracy of other nonlinear approaches, but with the significant advantage of learning the interpretable linear coefficients as well.
@@ Appendix A Equivalence of Adaptive Ridge regression and Lasso
The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact [Tibshirani, 1996, Grandvalet, 1998, Grandvalet and Canu, 1998]. Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
@@
Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and $\bm{y}\in\mathbb{R}^{n}$ be the targets. In a linear model the relationship between target and data is assumed to be linear:
$\bm{y}=X\bm{\beta}+\epsilon\;,$ (8)
@@
$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|_{2}^{2}+\lambda\sum_{j=1}^{d}\beta_{j}^{2}\;,$ (9)
@@
$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}\;,$ (10)
@@
$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}}\;,$ (11)
@@
We can now use the constraint in Eq. (10):
@@
If we now substitute $\nu_{k}|\beta_{k}|=\mu^{1/2}$ in the Adaptive Ridge objective in Eq. (10), we get the Lasso objective:
$f(\bm{\beta},\bm{\nu})$ $=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}$ (14)
$=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}|\beta_{i}|\cdot|\beta_{i}|$
$=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\mu^{1/2}|\beta_{i}|$
$=\|X\bm{\beta}-\bm{y}\|^{2}+\mu^{1/2}\sum_{i=1}^{d}|\beta_{i}|.$
which gives the relationship with the lasso turns out to be a Lasso regression, with the Lasso penalty parameter at optimum to be $\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$. ∎
Appendix B Training through profile likelihood
The proposed model in Eq. (2) has two sets of parameters: the linear coefficients $\bm{\beta}$ and the neural network parameters $\bm{u}$. Since the neural network learn the nonlinear corrections over the linear predictions, the two sets of parameters are clearly entangled and in practice optimizing them both with gradient descent is tricky. Furthermore, to increase the interpretability of the model we also encourage sparse coefficients and a sparse first fc layer of the the neural network. In this section we show how to make the optimization more effective and, at the same time, how to encourage sparsity. In particular, we propose a novel optimization method based on profile likelihood with adaptive ridge regression. The profile likelihood allows for an analytical update of the linear coefficients while adaptive rigde regression provides sparsity as in Lasso regression. We further show that we can extend such an approach to generalized linear models and we showcase it in detail for logistic regression.
@@ B.1 Regression setting: adaptive ridge regression
Let $\mathcal{L}(\bm{\beta},\bm{u})$ be the joint likelihood in Eq. (3). First, we fix the neural network parameters and analytically solve the maximization problem to obtain $\hat{\bm{\beta}}$. Then, we substitute $\hat{\bm{\beta}}$ back into the joint likelihood, which results in the profile likelihood for $\bm{u}$: $\mathcal{L}(\hat{\bm{\beta}},\bm{u})$. We can then optimize $\bm{u}$ by minimizing the negative log-profile likelihood using gradient descent. We iterate over these steps until convergence. More specifically, we draw the inspiration from the Adaptive Ridge (AR) regression and exploit its equivalence to Lasso regression [Tibshirani, 1996, Grandvalet, 1998, Grandvalet and Canu, 1998].
We already showed that the proposed model in Eq. (2) can be re-written in matrix form as:
@@
$\bm{y}=B_{\bm{u}}\bm{\beta}+\epsilon,$ (16)
where $\bm{y}\in\mathbb{R}^{n}$ denotes the targets. Adaptive Ridge regression minimizes the following objective:
$\min_{\bm{\beta}}\|B_{\bm{u}}\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (17)
where $\nu_{j}>0$ are the penalty parameters for each coefficient $\beta_{j}$, and $\lambda>0$ is a predefined parameter. To avoid divergent solutions we use the re-parameterization employed in Grandvalet [1998], Grandvalet and Canu [1998]:
@@
$\min_{\bm{\gamma}}\|\bm{y}-B_{\bm{u}}D_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}c_{i}^{2}=d,\quad c_{i}>0,$ (19)
where $D_{\bm{c}}$ is a diagonal matrix whose diagonal elements are given by the vector $\bm{c}\coloneqq[c_{1},\ldots,c_{d}]^{T}$. With this re-parameterized problem, we can define the profile likelihood. If we fix $\bm{c}$ and $\bm{u}$, $\hat{\bm{\gamma}}$ can be analytically obtained in closed form:
$\hat{\bm{\gamma}}=\left(D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\bm{y},$ (20)
We can then get $\hat{\bm{\beta}}$ in closed form:
$\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}\;,$ (21)
where $\circ$” denoted element-wise multiplication. We can now substitute $\hat{\bm{\beta}}$ back into the original problem and take the constraint into consideration. This leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ and neural network parameters $\bm{u}$:
$\min_{\bm{c},\bm{u}}\|\bm{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}\|\bm{c}\|^{2},$ (22)
and we can optimize it through gradient descent. The training algorithm for this re-parameterized adaptive ridge regression is summarized in Alg. 1.
@@
0:  Training data $X\in\mathbb{R}^{n\times d}$, training target $\bm{y}\in\mathbb{R}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameter $\lambda$, learning rate $\eta$, maximum iterations $T$.
1:  Initialize network parameters $\bm{u}$ and scaling coefficients $\bm{c}$ (with $c_{i}>0$).
@@
5:     $\hat{\bm{\gamma}}=\bigl{(}D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\bigr{)}^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\bm{y}$ // Solve for $\hat{\bm{\gamma}}$
@@
7:     $L=\|\bm{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}\|\bm{c}\|^{2}$ // Compute loss function
@@ B.2 Logistic regression: Iteratively reweighted least squares
The procedure followed in the previous section can be extended to generalized linear models. Here we showcase how to do it for logistic regression. Logistic regression is a classification model that describes the probability that a binary outcome $y\in\{0,1\}$ occurs given a feature vector $\bm{x}\in\mathbb{R}^{d}$:
$P(y=1|\bm{x})=\sigma(\bm{x}^{T}\bm{\beta})$ (23)
@@
$\mathcal{L}(\bm{\beta})=-\sum_{i=1}^{n}[y_{i}\log\sigma(\bm{x}_{i}^{T}\bm{\beta})+(1-y_{i})\log(1-\sigma(\bm{x}_{i}^{T}\bm{\beta}))]+\sum_{j=1}^{d}\nu_{j}\beta_{j}^{2}\quad\text{ s. t. }\quad\sum_{j=1}^{d}\frac{1}{\nu_{j}}=\frac{d}{\lambda}$ (24)
Now we can use the same re-parameterization used in the previous section for regression. We get the following optimization problem:
$\begin{gathered}\min_{\bm{\gamma}}-\sum_{i=1}^{n}\left[y_{i}\log\sigma\left(\sum_{j=1}^{p}x_{ij}c_{j}\gamma_{j}\right)+(1-y_{i})\log\left(1-\sigma\left(\sum_{j=1}^{p}x_{ij}c_{j}\gamma_{j}\right)\right)\right]+\lambda\|\bm{\gamma}\|^{2}\\ \text{ s.t. }\quad\sum_{j=1}^{d}c_{j}^{2}=d,\quad c_{i}>0\end{gathered}$ (25)
With the reparametrized objective we can consider the profile likelihood by fixing $\bm{c}$ and update $\bm{\gamma}$. To update $\bm{\gamma}$ we use the iteratively reweighted least squares (IRLS) algorithm. The associated update reads as:
$\bm{\gamma}^{(t+1)}=(\tilde{X}^{T}W^{(t)}\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W^{(t)}\mathbf{z}^{(t)}$ (26)
where $\tilde{X}$ is the modified design matrix with entries $\tilde{x}_{ij}=c_{j}x_{ij}$, $W^{(t)}$ is a diagonal weight matrix with entries $w_{i}^{(t)}=p_{i}^{(t)}(1-p_{i}^{(t)}),p_{i}^{(t)}=\frac{1}{1+e^{-\tilde{\bm{x}}_{i}^{T}\bm{\gamma}^{(t)}}}$, and $\mathbf{z}^{(t)}$ is the the vector of working response with entries $z_{i}^{(t)}=\bm{x}^{T}\bm{\beta}+\frac{y_{i}-p_{i}^{(t)}}{w_{i}^{(t)}}$. Note that in our model, our design matrix $X$ contains the nonlinear term as well so we would need to replace $X$ with $B_{\bm{u}}=X+X\circ G_{\bm{u}}$. We report the detailed optimization algorithm for the logistic regression case in Alg. 2.
Algorithm 2 IRLS Training for Logistic Regression
0:  Training data $X\in\mathbb{R}^{n\times d}$, target labels $\bm{y}\in\{0,1\}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameters $\lambda$, learning rate $\eta$, maximum iterations $T$.
1:  Initialize parameters $\bm{u}$, scaling coefficients $\bm{c}$ (with $c_{i}>0$), and regression coefficients $\bm{\beta}$.
2:  for $t=1,\dots,T$ do
3:     $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ // Compute transformed design matrix
4:     $D_{\bm{c}}=\operatorname{diag}(\bm{c})$ // Form diagonal scaling matrix
5:     $\tilde{X}=B_{\bm{u}}D_{\bm{c}}$
6:     $\hat{\bm{y}}=B_{\bm{u}}\bm{\beta}$ // Compute logits
7:     $\bm{\pi}=\sigma(\hat{\bm{y}})$
8:     $W=\operatorname{diag}(\bm{\pi}\circ(1-\bm{\pi}))$
9:     $\mathbf{z}=\hat{\bm{y}}+(\bm{y}-\bm{\pi})/\operatorname{diag}(W)$
10:     $\bm{\gamma}=(\tilde{X}^{T}W\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W\mathbf{z}$
11:     $\bm{\beta}=\bm{c}\circ\hat{\bm{\gamma}}$ // Update $\bm{\beta}$
12:     $L=\operatorname{BCE}(\hat{\bm{y}},\bm{y})+\lambda\sum_{i}c_{i}^{2}$ // Compute total loss
13:     $\bm{c}\leftarrow\bm{c}-\eta\nabla_{\bm{c}}L,\quad\bm{u}\leftarrow\bm{u}-\eta\nabla_{\bm{u}}L$ // Gradient descent step
14:  end for
15:  Return $\bm{\beta},\bm{c},\bm{u}$

### v2

@@ 5 Conclusions
In this paper, we propose a nonlinear interpretable model (NIMO). Unlike other hybrid approaches, our model preserves the interpretability of linear coefficients, capturing the additive contribution of each feature to the prediction, while also accommodating nonlinear interactions. Interpretability is formalized through marginal effects, with the marginal effects at the mean (MEM) providing global, population-level summaries that align directly with the linear coefficients. To enable effective training, we introduce an optimization algorithm based on parameter elimination, and enforce sparsity through adaptive ridge regularization. We demonstrate the effectiveness of the proposed approach on both synthetic and real datasets, showing competitive predictive performance while delivering meaningful and globally consistent interpretations. A key limitation of the approach is that it assumes the data can be described as a linear model with nonlinearity corrections; highly nonlinear datasets or features with self-interactions (e.g., $x_{j}f(x_{j})$) may not be fully captured. In such cases, careful feature engineering may be required to preserve interpretability.
@@ Appendix A Equivalence of Adaptive Ridge regression and Lasso
The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact (Tibshirani, 1996; Grandvalet, 1998; Grandvalet & Canu, 1998). Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
@@
Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and $\mathbf{y}\in\mathbb{R}^{n}$ be the targets. In a linear model the relationship between target and data is assumed to be linear:
$\mathbf{y}=X\bm{\beta}+\epsilon\;,$ (9)
@@
$\min_{\bm{\beta}}\|X\bm{\beta}-\mathbf{y}\|_{2}^{2}+\lambda\sum_{j=1}^{d}\beta_{j}^{2}\;,$ (10)
@@
$\min_{\bm{\beta}\,\nu}\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}\;,$ (11)
@@
$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}}\;,$ (12)
@@
If we now substitute $\nu_{k}|\beta_{k}|=\mu^{1/2}$ in the Adaptive Ridge objective in Eq. 11, we get the Lasso objective:
$f(\bm{\beta},\bm{\nu})$ $=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}$ (14)
$=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}|\beta_{i}|\cdot|\beta_{i}|$
$=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\mu^{1/2}|\beta_{i}|$
$=\|X\bm{\beta}-\mathbf{y}\|^{2}+\mu^{1/2}\sum_{i=1}^{d}|\beta_{i}|.$
which gives the relationship between adaptive ridge regression and Lasso regression, with the Lasso penalty parameter at optimum to be $\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$:
@@
Appendix B Training via parameter elimination
The proposed model in Eq. 2 has two sets of parameters: the linear coefficients $\bm{\beta}$ and the neural network parameters $\bm{u}$. Since these two sets of parameters are tightly entangled, jointly optimizing them via gradient descent is not straightforward in practice. Furthermore, to increase the interpretability of the model, we also aim to achieve sparsity in the learned coefficients. In this section, we show how to achieve these optimization goals via parameter elimination. Specifically, we eliminate $\bm{\beta}$ by expressing it in terms of the neural network parameters $\bm{u}$ and substituting it back into the optimization objective. The resulting problem can then be optimized effectively and efficiently using gradient descent. Sparsity is achieved by reformulating the problem as an adaptive ridge regression. We further show that this approach can be extended to generalized linear models, and we illustrate it in detail for logistic regression.
@@ B.1 Regression setting: adaptive ridge regression
We already showed that the proposed model in Eq. 2 can be re-written in matrix form as:
@@
$\mathbf{y}=B_{\bm{u}}\bm{\beta}+\epsilon,$ (17)
where $\mathbf{y}\in\mathbb{R}^{n}$ denotes the targets. Adaptive Ridge regression minimizes the following objective:
$\min_{\bm{\beta},\bm{u}}\|B_{\bm{u}}\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (18)
where $\nu_{i}>0$ are the penalty parameters for each coefficient $\beta_{i}$, and $\lambda>0$ is a predefined parameter. To avoid divergent solutions we use the re-parameterization employed in Grandvalet (1998); Grandvalet & Canu (1998):
@@
$\min_{\bm{\gamma},\bm{c},\bm{u}}\|\mathbf{y}-B_{\bm{u}}D_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}c_{i}^{2}=d,\quad c_{i}\geq 0,$ (20)
where $D_{\bm{c}}$ is a diagonal matrix whose diagonal elements are given by the vector $\bm{c}\coloneqq[c_{1},\ldots,c_{d}]^{T}$. Applying the parameter elimination trick, if we fix $\bm{c}$ and $\bm{u}$, $\bm{\gamma}$ can be expressed in a closed-form in terms of $\bm{c}$ and $\bm{u}$:
$\hat{\bm{\gamma}}(\bm{c},\bm{u})=\left(D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\mathbf{y},$ (21)
We can now substitute $\hat{\bm{\gamma}}(\bm{c},\bm{u})$ back into the original problem and take the constraint into consideration. This leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ and neural network parameters $\bm{u}$:
$\min_{\bm{c},\bm{u}}\|\mathbf{y}-B_{\bm{u}}D_{\bm{c}}\hat{\bm{\gamma}}(\bm{c},\bm{u})\|^{2}+\lambda\|\hat{\bm{\gamma}}(\bm{c},\bm{u})\|^{2}+\mu\|\bm{c}\|^{2},$ (22)
where $\mu$ is the Lagrangian multiplier introduced for $\bm{c}$, and we can optimize $\bm{c},\bm{u}$ together through gradient descent. The training algorithm for this re-parameterized adaptive ridge regression is summarized in Alg. 1.
@@
0: Training data $X\in\mathbb{R}^{n\times d}$, training target $\mathbf{y}\in\mathbb{R}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameter $\lambda$ and $\mu$, learning rate $\eta$, maximum iterations $T$.
1: Initialize network parameters $\bm{u}$ and scaling coefficients $\bm{c}$ (with $c_{i}\geq 0$).
@@
5:  $\hat{\bm{\gamma}}=\bigl(D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\bigr)^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\mathbf{y}$ // Solve for $\hat{\bm{\gamma}}$
@@
7:  $L=\|\mathbf{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\lambda\|\hat{\bm{\gamma}}\|+\mu\|\bm{c}\|^{2}$ // Compute loss function
@@ B.2 Logistic regression: Iteratively reweighted least squares
The procedure followed in the previous section can be extended to generalized linear models. Here we showcase how to do it for logistic regression. Logistic regression is a classification model that describes the probability that a binary outcome $y\in\{0,1\}$ occurs given a feature vector $\mathbf{x}\in\mathbb{R}^{d}$:
$P(y=1|\mathbf{x})=\sigma(\mathbf{x}^{T}\bm{\beta})$ (23)
@@
$\mathcal{L}(\bm{\beta})=-\sum_{i=1}^{n}[y_{i}\log\sigma(\mathbf{x}_{i}^{T}\bm{\beta})+(1-y_{i})\log(1-\sigma(\mathbf{x}_{i}^{T}\bm{\beta}))]+\sum_{j=1}^{d}\nu_{j}\beta_{j}^{2}\quad\text{ s. t. }\quad\sum_{j=1}^{d}\frac{1}{\nu_{j}}=\frac{d}{\lambda}$ (24)
Now we can use the same re-parameterization used in the previous section for regression. We get the following standard ridge-penalized logistic regression:
$\begin{gathered}\min_{\bm{\gamma},\bm{c}}-\sum_{i=1}^{n}\left[y_{i}\log\sigma\left((\mathbf{x}_{i}\odot\bm{c})^{T}\bm{\gamma}\right)+(1-y_{i})\log\left(1-\sigma\left((\mathbf{x}_{i}\odot\bm{c})^{T}\bm{\gamma}\right)\right)\right]+\lambda\|\bm{\gamma}\|^{2}\\ \text{ s.t. }\quad\sum_{j=1}^{d}c_{j}^{2}=d,\quad c_{j}\geq 0,\end{gathered}$ (25)
where $\bm{c}:=[c_{1},...,c_{d}]^{T}$, $\odot$ is element-wise multiplication. The same as what we done for regression, we can eliminate $\bm{\gamma}$ by expressing it in a closed-form with respect to $\bm{c}$. To get this closed-form expression of $\bm{\gamma}$, we use the iteratively reweighted least squares (IRLS) algorithm. The associated update reads as:
$\bm{\gamma}^{(t+1)}(\bm{c})=(\tilde{X}^{T}W^{(t)}\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W^{(t)}\mathbf{z}^{(t)}$ (26)
where $\tilde{X}$ is the modified design matrix with entries $\tilde{x}_{ij}=c_{j}x_{ij}$, $W^{(t)}$ is a diagonal weight matrix with entries $w_{i}^{(t)}=p_{i}^{(t)}(1-p_{i}^{(t)}),p_{i}^{(t)}=\frac{1}{1+e^{-\tilde{\bm{x}}_{i}^{T}\bm{\gamma}^{(t)}}}$, and $\mathbf{z}^{(t)}$ is the the vector of working response with entries $z_{i}^{(t)}=\tilde{\bm{x}}_{i}^{T}\bm{\gamma}^{(t)}+\frac{y_{i}-p_{i}^{(t)}}{w_{i}^{(t)}}$. We then substitute $\bm{\gamma}^{(t+1)}(\bm{c})$ back into the original problem, resulting the optimization over $\bm{c}$ only. Note that in our model, our design matrix $X$ contains the nonlinear term as well, so we would need to replace $X$ with $B_{\bm{u}}=X+X\circ G_{\bm{u}}$. The rest goes the same as the case for regression.